Tool Manufacturing has an expected EBIT of $83,000 in perpetuity and a tax rate of 25 percent. The company has $145,000 in outstanding debt at an interest rate of 6.5 percent and its unlevered cost of capital is 14 percent.

  

What is the value of the company according to MM Proposition I with taxes?
Answer #1

Value of unlevered company = EBIT*(1-t)/unlevered cost of capital

Value of unlevered company = 83000*(1-0.25)/0.14 = 444642.86

Value of levered company = Value of unlevered company + debt*(tax rate) = 444642.86 + 145000*(0.25) = 480892.86

Answer : $480892.86 (Thumbs up please)